To make a certain household paint, you need to mix 3/4L of white paint, 15ml of red paint and 10ml of yellow paint. If you need a total volume of 31 litres of this paint, how much of each colour must you mix?


1
Expert's answer
2022-02-14T16:16:56-0500
"\dfrac{3}{4}(1000\ mL)=750\ mL""750+15+10=775(mL)"

Divide by 5


"150\ parts+3\ parts+2\ parts=155\ parts"

The paint mixture as a ratio is "150:3:2."


"31\ L" corresponds to "155\ parts"

"0.2\ L" corresponds to "1\ part"

You need to mix 30 L of white paint with 0.6 L of red paint and 2 L yellow paint to obtain 31 litres of this paint